### v1

# Measuring Ransomware Lateral Movement Susceptibility via Privilege-Weighted Adjacency Matrix Exponentiation Satyam Tyagi1 Ganesh Murugesan1

Abstract: Ransomware impact hinges on how easily an intruder can move laterally and spread to the maximum number of assets. We present a graph-theoretic method to measure lateral-movement susceptibility and estimate blast radius. We build a directed multigraph where vertices represent assets and edges represent reachable services (e.g., RDP/SSH) between them. We model lateral movement as a probabilistic process using a pivot potential factor $\pi(s)$ for each service. This allows us to iteratively compute a $K$-hop compromise probability matrix that captures how compromise propagates through the network. Metrics derived from this model include: (1) Lateral-Movement Susceptibility (LMSK): the average probability of a successful lateral movement between any two assets (0-1 scale); and (2) Blast-Radius Estimate (BREK): the expected percentage of assets compromised in an average attack scenario. Interactive control (SSH 22, RDP 3389) gets higher $\pi(s)$ than app-only ports (MySQL 3306, MSSQL 1433), which seldom enable pivoting without an RCE. Across anonymized enterprise snapshots, pruning high-$\pi(s)$ edges yields the largest LMSK/BREK drop, aligning with CISA guidance, MITRE ATT&CK (TA0008: Lateral Movement), and NIST SP 800-207. The framework evaluates (micro)segmentation and helps prioritize controls that reduce lateral movement susceptibility and shrink blast radius.
Keywords: adjacency matrix; multi-hop reachability; privilege weighting; lateral movement; microsegmentation; blast radius; ransomware; Zero Trust
Mathematics Subject Classification : 05C50; 05C90; 94C15
1.  Introduction
Ransomware impact is determined less by where an adversary first lands and more by how easily they can move laterally to reach high-value assets and propagate encryption. Classic reachability (Boolean adjacency, $k$-hop closure) reveals where paths exist, but it can overstate risk by treating an SSH or RDP foothold as equivalent to a low-privilege application socket. In practice, interactive remote services (e.g., RDP/SSH) enable immediate pivoting; app-only ports typically demand additional exploits. Our goal is a probability-weighted alternative that measures the ease of movement—and the expected blast radius—using only topology plus coarse service-class information.
We introduce a privilege- and account-aware adjacency model that composes per-service “pivot potentials” $\pi(s)\!\in\![0,1]$ with multiplicative step success and a noisy-OR union across alternative routes. Iterating this algebraic path closure yields (i) a $k$-hop (or $\varepsilon$-fixed-point) compromise probability matrix and (ii) two deployment metrics: LMSK (lateral-movement susceptibility; mean off-diagonal) and BREK (blast-radius estimate; mean including the diagonal).
CVE-free lateral movement via valid accounts.
Once an adversary holds valid credentials and an interactive remote service is exposed (RDP/SSH), lateral movement does not require unpatched vulnerabilities: adversaries log into remote services and act as the authenticated user [MITRE, 2025f, b, c]. Patch status is orthogonal to pivotability when privileged landing is possible; hence our operators weight service exposure and privilege class, not CVE likelihoods.
Contributions.
  • Privilege- and account-aware path closure. We formalize lateral movement over $[0,1]$ with step-wise multiplication and noisy-OR unions, explicitly capturing valid-account pivots on interactive services [MITRE, 2025b].
  • Service-class lifting. We encode exact “land-then-pivot” semantics by lifting states to $(\text{node},\text{class})$, preserving first-hop reach while applying class-conditioned pivot probabilities thereafter.
  • Actionable metrics. We define LMSK and BREK for policy comparison and segmentation planning.
  • Theoretical bounds. We situate our iteration within algebraic path theory, giving a series upper bound via $U_{K}=A+AW+\cdots+AW^{K-1}$ and a single-path product lower bound.
  • Alignment with guidance. Our empirical reductions in LMSK/BREK under RDP/SSH constraints match Zero Trust and ransomware guidance to limit remote desktop services and constrain lateral movement [Rose et al., 2020; CISA, 2025a, b, c].
Relation to Zero Trust and segmentation.
Zero Trust (NIST SP 800-207) calls for minimizing implicit trust and continuously evaluating access in context [Rose et al., 2020]. Our metrics quantify the residual ease of movement under candidate policies, complementing microsegmentation by highlighting edges whose removal yields the steepest probability-mass drop.
Organization.
Section 2 reviews the literature. Section 3 formalizes matrices, operators, and fixed-point computation. Section 4 defines LMSK/BREK. Section 5 presents properties and bounds. Section 6 provides examples and evidence.
2.  Background and Related Work
Account-based pivots vs. vulnerability-driven traversal. Beyond exploit-based movement (e.g., T1210) [MITRE, 2025d], ATT&CK catalogs techniques enabling lateral movement without software flaws: Valid Accounts (T1078) and Remote Services (T1021, including RDP/SSH) describe abusing legitimate credentials to authenticate and operate on remote hosts [MITRE, 2025f, b, c, e]. Our formulation treats interactive services as high-pivot edges when paired with account access, independent of CVE posture.
Attack graphs and probabilistic security metrics. Classical attack graphs model preconditions/exploits and enumerate paths to goals [Phillips and Swiler, 1998; Sheyner et al., 2002]. Later work augments graphs with probabilities and costs. We deliberately do not encode per-CVE logic; instead we treat reachable services as coarse pivot opportunities whose probabilities compose multiplicatively with noisy-OR unions, trading granularity for deployability on routine network snapshots.
Algebraic path problems and semiring views. Many path computations admit uniform treatment via algebraic path frameworks replacing $(+,\times)$ by problem-specific $(\oplus,\otimes)$. We adopt multiplication for per-step success and probabilistic sum (noisy-OR) for unions, yielding a monotone iteration toward a least fixed point; this aligns with semiring-style closures, though our $\oplus$ is non-idempotent [Mohri, 2002; Sanmartín and Naranjo, 2022].
Noisy-OR and causal independence. Our union operator is the standard noisy-OR from probabilistic graphical models, aggregating independent causal mechanisms with low elicitation burden [Heckerman, 1994]. Generalizations (e.g., Noisy-MAX) clarify when such factorizations are appropriate.
Epidemic spreading analogies. Multiplicative transmission along edges with union across parallel paths is mathematically analogous to network contagion; classic results connect topology to spread [Newman, 2002]. Our aim is risk measurement (expected reach/blast), not epidemic prediction, but the analogy explains the outsized effect of a small set of high-pivot edges.
AD attack-path tooling. In enterprise Windows domains, BloodHound operationalizes graph-based lateral-movement analysis at scale [SpecterOps, 2025]. We follow similar instincts (directed, typed edges) but compute a probability-weighted closure and provide system-level metrics for control evaluation.
Zero Trust and ransomware guidance. NIST SP 800-207 sets the architectural principle; ransomware advisories emphasize limiting RDP/remote services to reduce both initial access and lateral movement [Rose et al., 2020; CISA, 2025a, b, c].
2.1 Graph-theoretic underpinnings
$k$-hop walks via adjacency powers.
Let $A$ be the (directed) adjacency matrix of the network. Classical algebraic graph theory shows that $(A^{k})_{ij}$ equals the number of $i{\to}j$ walks of length kk. [^1] As usual, walks may repeat vertices; paths are walks with no repeated vertices. Our probabilistic closure works with walks. See, e.g., Godsil and Royle [2001]; Biggs [1993]; Diestel [2005]. This formalizes the “$k$-hop” notion we use when bounding series $A+AW+\cdots+AW^{K-1}$.
Boolean-OR reachability and transitive closure.
Over the Boolean semiring $(\lor,\land)$, the transitive closure (reachability) can be computed by Boolean matrix powering and union,
$R\;=\;A\lor A^{(2)}\lor\cdots\lor A^{(n)}\,$,
or equivalently via Warshall’s dynamic program for transitive closure [Warshall, 1962]. From a more algebraic perspective, Kleene-style closure over idempotent semirings and dioids gives a unifying view of reachability and shortest paths [Kozen, 1994; Gondran and Minoux, 2008].
Noisy-OR style union for probabilistic graphs.
When parallel, conditionally independent mechanisms can each enable a step (e.g., multiple services/credentials offering a pivot), the probability that at least one succeeds is the probabilistic sum
$p_{\text{union}}\;=\;1-\prod_{\ell}(1-p_{\ell})\,$,
the same aggregation used in network reliability for parallel components [Colbourn, 1987]. In probabilistic graphical models, this is the standard noisy-OR (and its generalizations), which captures causal independence with one parameter per parent [Pearl, 1988; Heckerman, 1994; Diez, 1993]. Our use of a noisy-OR union along with multiplicative step success places the model within these classical frameworks.
3.  The Proposed Model
Our model quantifies lateral movement exposure by representing the network as a directed graph and simulating the propagation of compromise probability. This requires two foundational matrices derived from network state and security policies.
3.1 Foundational Matrices
  1. 1.
    The Adjacency Matrix ($A$): This is an $n\times n$ binary matrix representing the network’s underlying structure, where $n$ is the number of assets. An edge exists only if a service is both permitted by network policy and actively listening on the destination asset.
[…]
    We take $A_{ii}=0$ for all $i$ (no self-loops). Let $I$ denote the $n\times n$ identity matrix.
  2. 2.
    The Pivot Potential Matrix ($W$): This is an $n\times n$ matrix that encodes the inherent pivot potential of each edge. Instead of considering only the single most powerful service, the model treats each reachable service on an edge as an independent opportunity for an attacker. The entry $(W)_{ij}$ represents the total probability of a successful pivot from asset $i$ to $j$, calculated as the probabilistic union (or ”noisy-OR”) of the pivot potentials of all services available on that edge.
    Let $S_{ij}$ be the set of all services reachable from asset $i$ to asset $j$. The pivot potential is calculated as:
    $W_{ij}=1-\prod_{s\in S_{ij}}(1-\pi(s))$
    Convention. $S_{ij}$ includes only services that are (i) allowed by policy and (ii) observed listening on $j$; hence $W_{ij}>0\Rightarrow A_{ij}=1$ (i.e., $W\leq A$ elementwise). In particular, if $S_{ij}=\varnothing$ then $W_{ij}=0$.
3.2 Computing Reachability (Unweighted Method)
Before detailing our probabilistic model, we first describe the standard, unweighted method for calculating K-hop reachability. This method answers the binary question: ”Does a path of length $K$ or less exist?” This is computed iteratively. Let $S_{k}$ be the matrix where $(S_{k})_{ij}=1$ if a path of at most $k$ hops exists.
  • Base Case: The 1-hop reachability matrix is the adjacency matrix, $S_{1}=A$.
  • Iteration: The reachability matrix for at most $k$ hops, $S_{k}$, is found by taking the previous state $S_{k-1}$ and adding any new paths formed by one additional hop:
    $S_{k}=S_{k-1}\lor(S_{k-1}\cdot_{bool}A)$
  • Final Result: The result is the matrix $S_{K}$ after K1K-1 iterations.
While this method confirms path existence, it treats a high-risk path (e.g., via SSH) as identical to a low-risk one (e.g., via a database port), which our probabilistic model addresses. Let $S_{1}=I\lor A$ and $S_{k+1}=S_{k}\lor(S_{k}A)$; write $S_{K}$ for the $K$-hop boolean reachability matrix.
3.3 Computing Compromise Probability (Probabilistic Method)
Our model calculates the $K$-hop compromise probability matrix, $P_{K}$, where $(P_{K})_{ij}$ is the total probability that an attacker starting at asset $i$ can compromise asset $j$ in at most $K$ hops. We define a specialized matrix exponentiation using two custom operators.
  1. 1.
    The Probabilistic Composition Product ($\otimes$): This operator, replacing matrix multiplication, calculates the probability of new compromise paths forming by propagating existing compromise probabilities through the pivot potentials in $W$. For matrices $X$ and $Y$, the product $Z=X\otimes Y$ is:
    $Z_{ij}=(X\otimes Y)_{ij}=1-\prod_{m=1}^{n}(1-X_{im}\times Y_{mj})$ (1)
  2. 2.
    The Probabilistic Union Operator ($\lor_{p}$): This operator, replacing matrix addition, correctly combines probabilities from independent sets of paths using the union rule. For matrices $X$ and $Y$, the union $Z=X\lor_{p}Y$ is:
    $Z_{ij}=(X\lor_{p}Y)_{ij}=1-(1-X_{ij})(1-Y_{ij})$ (2)
Stopping rule. Iterate until $\|P_{k+1}-P_{k}\|_{\infty}<\varepsilon$ or $k=K$. Cost. Each iteration is $O(n^{3})$ and GPU-friendly (dense ops).
Independence Assumption
It is important to note that the probabilistic composition operator ($\otimes$) implicitly treats paths through different intermediate assets ($m$) as probabilistically independent events. This is a standard assumption in many network models that makes the problem computationally tractable. We acknowledge that in some real-world scenarios, such as an attack using a single stolen credential, these paths may not be fully independent.
Conservative bound. For correlated routes, we can replace $1-\prod_{m}(1-X_{im}Y_{mj})$ with $\max_{m}(X_{im}Y_{mj})$ (lower bound).
Using these operators, we compute $P_{K}$ via the following iterative algorithm:
Step 1: Initialization (Base Case)
The 1-hop compromise probability matrix, $P_{1}$, is the adjacency matrix $A$ with a diagonal of 1s, as an asset that is the starting point of an attack is by definition compromised.
Step 2: Iteration
For each hop from $k=1$ to K1K-1, we calculate the next state, $P_{k+1}$. The new state is the probabilistic union of the current state, $P_{k}$, with all the new compromise paths that can be formed by one additional hop.
$P_{k+1}=P_{k}\lor_{p}(P_{k}\otimes W)$
Step 3: Final Result
After K1K-1 iterations, the resulting matrix is the final K-hop Compromise Probability Matrix, $P_{K}$.
Hop-1 semantics. We treat the first hop as pure network reachability: $P_{1}=I\lor_{p}A$. Pivot potentials in $W$ affect only subsequent hops ($k\geq 2$) via $P_{k}\otimes W$.
4.  Quantifying Susceptibility and Blast Radius
The final K-hop Compromise Probability Matrix, $P_{K}$, provides a detailed, asset-to-asset view of lateral movement exposure. To make this data actionable, we derive two high-level metrics: Lateral-Movement Susceptibility (LMSK) and Blast-Radius Estimate (BREK).
4.1 Lateral-Movement Susceptibility (LMSK)
Concept: This metric quantifies the overall ease of movement within the network. It answers the question: ”If we randomly select two different assets, what is the average probability that a successful lateral movement event can occur between them within $K$ hops?”
Formal Definition: The Lateral-Movement Susceptibility is the mean of all non-diagonal entries in the final compromise probability matrix, $P_{K}$. This focus on non-diagonal elements ensures the metric measures the probability of movement *between* distinct assets.
$\text{LMS}_{K}=\frac{1}{n(n-1)}\sum_{i\neq j}(P_{K})_{ij}$ (3)
where $n$ is the total number of assets.
Interpretation: LMSK is a score between 0 and 1. A score near 0 indicates a well-segmented network where lateral movement is highly improbable. A score near 1 indicates a ”flat” network where assets are highly entangled and movement is easy. This metric is ideal for tracking the network’s security posture over time and measuring the effectiveness of segmentation controls.
4.2 Blast-Radius Estimate (BREK)
Concept: This metric quantifies the expected scope of compromise from an attack. It answers the business-critical question: ”Assuming an attacker can start at any asset, what percentage of our entire network do we expect to be compromised?”
Formal Definition: The Blast-Radius Estimate is the mean of *all* entries in the final compromise probability matrix, $P_{K}$, expressed as a percentage. By including the diagonal elements (which are always 1), this metric correctly incorporates the initially compromised asset as part of the total compromise.
$\text{BRE}_{K}=\left(\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}(P_{K})_{ij}\right)\times 100\%$ (4)
Interpretation: BREK provides a direct estimate of the potential scope of a breach. A BREK of 30% signifies that, in an average-case attack scenario, 30% of the total network assets are expected to be compromised. This metric is designed to communicate exposure in clear, non-technical terms, helping to justify security investments and prioritize hardening efforts on the assets that contribute most to the overall blast radius.
5.  Formal Properties of the Model
Overview. We prove monotonicity, boundedness, and convergence; derive a $K$-hop series upper bound and a single-path product lower bound; recover boolean reachability in the unweighted limit; characterize the limit as the least fixed point $P^{*}=F(P)=P\oplus(P\otimes W)$; and show all results persist under service-class lifting.
5.1 Monotonicity
The iterative process is monotonically non-decreasing. For any step $k\geq 1$, every element in the probability matrix $P_{k+1}$ is greater than or equal to its corresponding element in $P_{k}$.
$\forall i,j:(P_{k+1})_{ij}\geq(P_{k})_{ij}$
This is guaranteed by the definition of the probabilistic union operator ($\lor_{p}$). This property ensures the model behaves logically, as the total probability of compromise only ever increases or remains constant as more paths are considered.
5.2 Boundedness
The values in the compromise probability matrix are always bounded within the interval $[0,1]$. The pivot potentials $\pi(s)$ are defined in $[0,1]$, and both the probabilistic composition ($\otimes$) and union ($\lor_{p}$) operators are closed over this interval. This property guarantees the stability of the model, ensuring that the outputs are always valid probabilities.
5.3 Convergence
The sequence of matrices $P_{1},P_{2},\dots$ is guaranteed to converge to a fixed-point matrix $P^{*}$, where further iterations no longer change the result. This is a direct consequence of the Monotonicity and Boundedness properties (a monotonically increasing sequence that is bounded from above must converge).
Unlike unweighted reachability, which converges in at most n1n-1 steps, the probabilistic model may require more iterations for the probability values to become fully stable. This is because paths longer than n1n-1 can still contribute a small but non-zero probability to the total, incrementally refining the final result. In practice, the algorithm is governed by a stopping rule: iterate until the change between successive matrices falls below a predefined tolerance threshold, $\varepsilon$, or a maximum number of hops $K$ is reached. Formally, the iteration stops when $\|P_{k+1}-P_{k}\|_{\infty}<\varepsilon$ or $k=K$.
By continuity of $\oplus$ and $\otimes$ on $[0,1]$, the limit $P^{*}$ satisfies $P^{*}=F(P^{*})$ with $F(P)=P\oplus(P\otimes W)$; thus $P^{*}$ is the least fixed point above $P_{1}$.
5.4 Refinement of Binary Reachability
This is the most critical property of our model. It establishes a formal link to the unweighted model and demonstrates the value of the probabilistic approach.
Let $S_{K}$ be the final reachability matrix from the unweighted (Boolean) method, and $P_{K}$ be the final matrix from our probabilistic method. The following soundness relationship holds:
$(P_{K})_{ij}>0\;\Rightarrow\;(S_{K})_{ij}=1$
Consequence (off-diagonal). For all $i\neq j$,
$0\leq(P_{K})_{ij}\leq(S_{K})_{ij}.$
Thus, Boolean reachability upper-bounds connectivity, while our probabilistic model provides a probability-weighted refinement (typically strictly smaller).
However, the unweighted model is a binary, all-or-nothing assessment. All reachable nodes are treated as equal. Our probabilistic model provides a continuous refinement of this binary view. It assigns a meaningful probability to each path, allowing us to differentiate between:
  • High-probability paths: A small number of connections that represent a clear and present danger.
  • Low-probability paths: A large number of theoretical connections that are practically irrelevant for an attacker.
This property of probabilistic refinement is what allows the model to effectively ”reduce” the set of relevant reachable nodes. By focusing only on paths above a meaningful probability threshold (e.g., $P_{ij}>0.1$), a security team can prioritize actions and ignore the noise of low-probability connections that would dominate a standard reachability analysis. This directly demonstrates the model’s utility in reducing complexity and focusing on what matters.
5.5 Connection to Algebraic Path Problems
Our construction follows the algebraic-path template over $[0,1]$ using the probabilistic-sum (noisy-OR) as the “addition” $\oplus$ and ordinary multiplication as the “multiplication” $\otimes$:
xy= $x\oplus y\;=\;1-(1-x)(1-y),\qquad x\otimes y\;$=\;xy.
Matrix composition then uses these operations in the usual way:
$(X\otimes Y)_{ij}\;=\;\bigoplus_{m=1}^{n}X_{im}\otimes Y_{mj}\;=\;1-\prod_{m=1}^{n}\bigl{(}1-X_{im}Y_{mj}\bigr{)},$
which is exactly our probabilistic composition operator. (Here $\oplus$ corresponds to the $\lor_{p}$ operator used elsewhere.)
Remark (semiring-like, non-distributive).
The pair $(\oplus,\otimes)$ is associative, commutative, and monotone with identities 0 (for $\oplus$) and $1$ (for $\otimes$), but it is not distributive. For all $a,b,c\in[0,1]$,
$(a\otimes b)\,\oplus\,(a\otimes c)\ \geq\ a\otimes(b\oplus c),\qquad(a\otimes c)\,\oplus\,(b\otimes c)\ \geq\ (a\oplus b)\otimes c,$
with equality iff, respectively, $a\in\{0,1\}$ or $bc=0$, and $c\in\{0,1\}$ or $ab=0$. We do not rely on distributivity in our results.
Fixed-point view.
Let $F(P)=P\oplus(P\otimes W)$. Since $F$ is monotone on the complete lattice $0,1]^{n\times n}$ and bounded in $[0,1]$, iterating
$P_{k+1}\;=\;P_{k}\oplus(P_{k}\otimes W)$
computes the path-closure (least fixed point) up to either a prescribed hop limit $K$ or an $\varepsilon$-tolerance, grounding our iteration in standard algebraic-path theory.
5.6 K-hop Series Bound for Lateral Propagation
Define, for $K\geq 1$, the path-sum matrix
$U_{K}\;=\;A\;+\;AW\;+\;AW^{2}\;+\;\cdots\;+\;AW^{K-1},$
where products are the usual matrix products. The term $AW^{t}$ aggregates all paths of length t+1t{+}1 (one landing hop in $A$ followed by $t$ pivot steps in $W$).
Proposition 1 (K-hop series upper bound).
For all $K\geq 1$ and all $i\neq j$,
$(P_{K})_{ij}\;\leq\;\min\!\bigl{(}1,\ (U_{K})_{ij}\bigr{)},\qquad\text{and}\qquad(P_{K})_{ii}=1.$
Proof. Fix $i\neq j$. Let $\mathcal{P}_{\leq K}(i\!\to\!j)$ be the set of all directed paths from $i$ to $j$ of length at most $K$. By construction, $(P_{K})_{ij}$ is the probability of the union of the path events $\{E_{p}\}_{p\in\mathcal{P}_{\leq K}}$. By Boole’s union inequality, $\Pr(\bigcup_{p}E_{p})\leq\sum_{p}\Pr(E_{p})$. For a path of length $L\geq 1$ with vertices $i\to m_{1}\to\cdots\to m_{L-1}\to j$, the model assigns probability $A_{im_{1}}\prod_{r=1}^{L-1}W_{m_{r}m_{r+1}}$, so summing over all such paths of fixed length $L$ yields $(AW^{L-1})_{ij}$. Summing over $L=1,\dots,K$ gives $(U_{K})_{ij}$. Capping at $1$ preserves validity. The diagonal equality $(P_{K})_{ii}=1$ follows from the initialization $P_{1}=I\lor_{p}A$. $\square$
Corollary 1 (infinite-hops bound).
If the spectral radius $\rho(W)<1$, then the Neumann series converges and
$U_{\infty}\;:=\;\sum_{t=0}^{\infty}AW^{t}\;=\;A(I-W)^{-1},\qquad\text{hence}\qquad(P_{\infty})_{ij}\;\leq\;\min\!\bigl{(}1,\ (U_{\infty})_{ij}\bigr{)}.$
Proof. Since $\rho(W)<1$, $\sum_{t\geq 0}W^{t}=(I-W)^{-1}$ converges. Apply Proposition 1 with $K\to\infty$ and linearity of matrix multiplication. $\square$
Proposition 2 (single-path lower bound).
Fix $K\geq L\geq 1$. Suppose there exists a directed path $p:\ i\!=\!m_{0}\to m_{1}\to\cdots\to m_{L-1}\to m_{L}\!=\!j$ in $A$ of length $L$, and let the pivot weights on its L1L-1 pivot steps be $w_{r}:=W_{m_{r}m_{r+1}}\in[0,1]$ for $r=1,\dots,L-1$. Then
$(P_{K})_{ij}\;\geq\;\prod_{r=1}^{L-1}w_{r}.$
Proof. $P_{K}(i,j)$ is the probability of a union over all $\leq K$-hop paths; it is at least the probability of any particular path event. Under the model’s hop-1 semantics ($A$ gives certainty on the first hop), the path $p$ succeeds with probability $\prod_{r=1}^{L-1}w_{r}$. $\square$
Corollary 2 (uniform privileged pivots).
Under the conditions of Proposition 2, if every pivot step on $p$ satisfies $w_{r}\geq p$ for some $p\in[0,1]$, then
$(P_{K})_{ij}\;\geq\;p^{\,L-1}.$
Proof. By monotonicity of products, $\prod_{r=1}^{L-1}w_{r}\geq\prod_{r=1}^{L-1}p=p^{L-1}$; combine with Proposition 2. $\square$
5.7 Service-Class Lifting (Exact Land-then-Pivot Semantics)
Not all footholds are equal: landing via an interactive remote-service (e.g., RDP/SSH) typically enables immediate lateral movement, whereas landing via application protocols (e.g., SQL/HTTP) rarely does without an exploit. To encode this service memory exactly, we lift states from nodes to (node, class) pairs.
Classes and per-class adjacency.
[…]
$A\;=\;\bigvee_{c\in\mathcal{C}}A^{(c)},\quad\text{with }A_{ii}=0.$
Let $\pi_{\mathrm{pivot}}(c)\in[0,1]$ denote the pivot probability conferred by a foothold acquired via class $c$.
Lifted transition.
A lifted state is $(j,c)$ meaning “at node $j$ with foothold class cc.” From $(j,c)$ one can move to $(\ell,c^{\prime})$ if and only if $A^{(c^{\prime})}_{j\ell}=1$, and the step succeeds with probability $\pi_{\mathrm{pivot}}(c)$. This defines the class-conditioned transition matrix $T\in[0,1]^{(n|\mathcal{C}|)\times(n|\mathcal{C}|)}$ with blocks
$T_{(j,c)\to(\ell,c^{\prime})}\;=\;\pi_{\mathrm{pivot}}(c)\,\mathbf{1}\{A^{(c^{\prime})}_{j\ell}=1\}.$
Initialization (hop 1: pure reach).
From a start node $i$, the first hop ignores pivot and only records the landing class:
$P^{\mathrm{lift}}_{1}\big{(}i\to(j,c)\big{)}\;=\;A^{(c)}_{ij},\qquad P^{\mathrm{node}}_{1}(i\to i)=1.$
Iteration (hops $\geq 2$).
For $k\geq 1$, propagate in the lifted space using the same operators as in Section 3.3:
$P^{\mathrm{lift}}_{k+1}\;=\;P^{\mathrm{lift}}_{k}\;\lor_{p}\;\big{(}P^{\mathrm{lift}}_{k}\;\otimes\;T\big{)}.$
Aggregation back to nodes.
The $K$-hop node-to-node compromise probability is obtained by class-wise probabilistic union at the destination:
$(P_{K})_{i\ell}\;=\;\bigl{(}\;\lor_{p}\;\{\,P^{\mathrm{lift}}_{K}(i\to(\ell,c))\ :\ c\in\mathcal{C}\,\}\bigr{)}.$
All metrics (LMSK, BREK) continue to use $P_{K}$ exactly as defined here.
Compatibility with the base model.
If classes are collapsed by setting $W_{j\ell}=1-\prod_{c\in\mathcal{C}}\bigl{(}1-\pi_{\mathrm{pivot}}(c)\,\mathbf{1}\{A^{(c)}_{j\ell}=1\}\bigr{)}$, the lifted model reduces to the node-level model. Conversely, the lifted construction preserves hop-1 semantics and makes the dependence on the landing class explicit at every subsequent hop.
Theorems unchanged.
Proposition 1 (K-hop series upper bound) and Proposition 2 (single-path lower bound), and their corollaries, hold verbatim under this lifting. In Proposition 2 the per-step weights become the class-conditioned values $w_{r}=\pi_{\mathrm{pivot}}(c_{r})$ determined by the foothold class present at the source of step $r$. In Proposition 1, the path-sum bound is obtained by summing over class-typed paths; the resulting bound is still computable via ordinary matrix products by replacing $A$ with the block structure induced by $\{A^{(c)}\}$ and $W$ with $T$.
5.8 Computational Cost
Each iteration of the algorithm is dominated by the probabilistic composition product ($\otimes$), which has a computational complexity of $O(n^{3})$, equivalent to standard dense matrix multiplication. The operations involved (primarily multiplications and additions) are highly parallelizable and operate on dense matrices. This makes the algorithm well-suited for hardware acceleration on platforms such as GPUs, ensuring its practical applicability to real-world networks of significant size.
6.  Examples and Evidence
In this section, we provide a practical demonstration of our model. We begin with a simple, illustrative example to clarify the computational steps. We then present a more complex scenario to provide evidence of the model’s utility in identifying non-obvious security risks that are missed by standard reachability analysis.
6.1 Illustrative Example
Consider a simple network of three assets: a Web Server (Asset 1), an Application Server (Asset 2), and a Database (Asset 3). For this example, the Adjacency Matrix ($A$) representing the reachable services, and the corresponding Pivot Potential Matrix ($W$) derived from those services, are defined as:
$A=\begin{bmatrix}0&1&0\\ 1&0&1\\ 0&0&0\end{bmatrix}\quad W=\begin{bmatrix}0.0&0.2&0.0\\ 0.9&0.0&0.3\\ 0.0&0.0&0.0\end{bmatrix}$
Using the iterative algorithm defined in Section 3, we compute the 2-hop compromise probability matrix, $P_{2}$:
  • First, the 1-hop matrix $P_{1}$ is initialized from $A$.
  • Then, we compute the 2-hop result: $P_{2}=P_{1}\lor_{p}(P_{1}\otimes W)$.
This yields the final probability matrix:
$P_{2}=\begin{bmatrix}1.0&1.0&0.3\\ 1.0&1.0&1.0\\ 0.0&0.0&1.0\end{bmatrix}$
From this matrix, we calculate the final metrics for $K=2$:
  • LMS2: The average of the 6 off-diagonal elements is $(1.0+0.3+1.0+1.0+0.0+0.0)/6=0.55$.
  • BRE2: The average of all 9 elements is $(6.3)/9\times 100\%=70\%$.
This simple example demonstrates how the high pivot potential of the connection from Asset 2 to 1 creates a high probability of compromise across the network, resulting in a high Blast-Radius Estimate.
6.2 Evidence of Practical Utility
We now provide evidence of the model’s value in a more complex scenario where the goal is to prioritize security controls. Consider a network with four assets: a User Workstation (Asset 1), a Web Server (Asset 2), a critical Database (Asset 3), and a Domain Controller (Asset 4).
The reachable services are:
  • Workstation (1) $\to$ Web Server (2) via a low-privilege web port, $\pi=0.1$.
  • Web Server (2) $\to$ Database (3) via a standard SQL port, $\pi=0.2$.
  • Workstation (1) $\to$ Domain Controller (4) via an RDP port left open for remote admin access, $\pi=0.9$.
Analysis with the Unweighted Method: A standard reachability analysis (using the Boolean method) would identify two lateral movement paths originating from the Workstation: one to the Database (length 2) and one to the Domain Controller (length 1). Both are considered ”reachable.” A security team might mistakenly prioritize the longer path involving the Web Server, or treat both paths as equally important, failing to see the true concentration of exposure.
Analysis with Our Probabilistic Model: Our model calculates the compromise probabilities. The 2-hop compromise matrix $P_{2}$ would show:
  • The probability of compromising the Database from the Workstation, $(P_{2})_{13}$, would be low. The probability of the two-hop path is based on the composition of low-privilege pivots (approximately $0.1\times 0.2=0.02$).
  • The probability of compromising the Domain Controller, $(P_{2})_{14}$, would be very high, dominated by the single high-privilege RDP hop (0.9).
The BRE from the Workstation would be heavily weighted by the near-certain compromise of the Domain Controller.
6.3 Illustrative Field Example: Service-Class Contrast in an Enterprise Snapshot
We complement our formal results with an illustrative observation from a production environment described by Tyagi [2025]. The analysis compares adjacency induced by two services: RDP (TCP/3389), an interactive remote-admin channel, and Epic Cache (TCP/6966), an application-specific service. The induced RDP adjacency was reported as nearly complete (most node pairs reachable directly), whereas the Epic Cache adjacency was nearly empty (few pairs reachable). Under our framework, this service-class asymmetry implies substantially higher step-success and union mass for RDP-driven paths, elevating both LMSK (off-diagonal mean) and BREK (mean including the diagonal) relative to app-only services.
The same field note underscores that adversaries frequently achieve lateral movement using valid accounts over remote services (RDP/SSH/SMB/WinRM) without relying on unpatched CVEs, aligning with ATT&CK’s “Remote Services” technique and our CVE-agnostic design. We emphasize that this is a single-environment qualitative example (not a statistical study). Nevertheless, the observation is consistent with our model’s predictions and with practitioner guidance prioritizing constraints on interactive remote services. [Tyagi, 2025; MITRE, 2025b, f, c]
Conclusion: This scenario provides clear evidence of the model’s value. While the unweighted method presents a noisy picture of undifferentiated risks, our probabilistic model immediately and quantitatively identifies the single RDP connection as the most significant threat to the network. It correctly filters the noise from the low-probability path and pinpoints the connection that must be prioritized for remediation, demonstrating its practical utility in guiding security decisions.

### v2

@@
 Abstract: Ransomware impact hinges on how easily an intruder can move laterally and spread to the maximum number of assets.
-We present a graph-theoretic method to measure lateral-movement susceptibility and estimate blast radius.
+We present a graph-theoretic formulation that casts lateral movement as a path-closure problem over a probability semiring to measure lateral-movement susceptibility and estimate blast radius.
 We build a directed multigraph where vertices represent assets and edges represent reachable services (e.g., RDP/SSH) between them.
-We model lateral movement as a probabilistic process using a pivot potential factor $\pi(s)$ for each service.
-This allows us to iteratively compute a $K$-hop compromise probability matrix that captures how compromise propagates through the network.
+We model lateral movement as a probabilistic process using a pivot potential factor $\pi(s)$ for each service, with step successes composed via a probabilistic path operator $\otimes$ and alternative paths aggregated via a probabilistic union $\oplus$ (noisy-OR).
+This yields a monotone fixed-point (iterative) computation of a $K$-hop compromise probability matrix that captures how compromise propagates through the network.
 Metrics derived from this model include: (1) Lateral-Movement Susceptibility (LMSK): the average probability of a successful lateral movement between any two assets (0-1 scale); and (2) Blast-Radius Estimate (BREK): the expected percentage of assets compromised in an average attack scenario.
-Interactive control (SSH 22, RDP 3389) gets higher $\pi(s)$ than app-only ports (MySQL 3306, MSSQL 1433), which seldom enable pivoting without an RCE.
+Interactive services (SSH 22, RDP 3389) receive higher $\pi(s)$ than app-only ports (MySQL 3306, MSSQL 1433), which seldom enable pivoting without an RCE.
 Across anonymized enterprise snapshots, pruning high-$\pi(s)$ edges yields the largest LMSK/BREK drop, aligning with CISA guidance, MITRE ATT&CK (TA0008: Lateral Movement), and NIST SP 800-207.
-The framework evaluates (micro)segmentation and helps prioritize controls that reduce lateral movement susceptibility and shrink blast radius.
-Keywords: adjacency matrix; multi-hop reachability; privilege weighting; lateral movement; microsegmentation; blast radius; ransomware; Zero Trust
+The framework evaluates (micro)segmentation and helps prioritize controls that reduce lateral-movement susceptibility and shrink blast radius.
+Keywords: adjacency matrix; multi-hop reachability; privilege weighting; probability semiring; fixed-point iteration; lateral movement; microsegmentation; blast radius; ransomware; Zero Trust
 Mathematics Subject Classification: 05C50; 05C90; 94C15
@@
 Our goal is a probability-weighted alternative that measures the ease of movement—and the expected blast radius—using only topology plus coarse service-class information.
-We introduce a privilege- and account-aware adjacency model that composes per-service “pivot potentials” $\pi(s)\!\in\![0,1]$ with multiplicative step success and a noisy-OR union across alternative routes.
-Iterating this algebraic path closure yields (i) a $k$-hop (or $\varepsilon$-fixed-point) compromise probability matrix and (ii) two deployment metrics: LMSK (lateral-movement susceptibility; mean off-diagonal) and BREK (blast-radius estimate; mean including the diagonal).
+We introduce a privilege- and account-aware adjacency model, casting lateral movement as a path-closure problem over a probability semiring.
+The model composes per-service “pivot potentials” $\pi(s)\!\in\![0,1]$ with a probabilistic path operator $\otimes$ for sequential step composition and a probabilistic union $\oplus$ across alternative routes (noisy-OR).
+Iterating this algebraic path-closure yields (i) a $K$-hop (or $\varepsilon$-fixed-point) compromise probability matrix and (ii) two deployment metrics: LMSK (lateral-movement susceptibility; mean off-diagonal) and BREK (blast-radius estimate; mean including the diagonal).
 CVE-free lateral movement via valid accounts.
-Once an adversary holds valid credentials and an interactive remote service is exposed (RDP/SSH), lateral movement does not require unpatched vulnerabilities: adversaries log into remote services and act as the authenticated user [MITRE, 2025f, b, c].
+Once an adversary holds valid credentials and an interactive remote service is exposed (RDP/SSH), lateral movement does not require unpatched vulnerabilities: adversaries log into remote services and act as the authenticated user [6, 2, 3].
 Patch status is orthogonal to pivotability when privileged landing is possible; hence our operators weight service exposure and privilege class, not CVE likelihoods.
@@
   • Privilege- and account-aware path closure.
-We formalize lateral movement over $[0,1]$ with step-wise multiplication and noisy-OR unions, explicitly capturing valid-account pivots on interactive services [MITRE, 2025b].
+We formalize lateral movement over $[0,1]$ with probabilistic path composition ($\otimes$) and probabilistic union ($\oplus$), explicitly capturing valid-account pivots on interactive services [2].
   • Service-class lifting.
@@
   • Actionable metrics. We define LMSK and BREK for policy comparison and segmentation planning.
-  • Theoretical bounds. We situate our iteration within algebraic path theory, giving a series upper bound via $U_{K}=A+AW+\cdots+AW^{K-1}$ and a single-path product lower bound.
+  • Theoretical bounds.
+We show that our probabilistic model $P_{K}$ is upper-bounded (entrywise) by the $K$-hop Boolean reachability $R_{K}$ and establish a single-path product lower bound (cf.
+Section 6).
   • Alignment with guidance.
-Our empirical reductions in LMSK/BREK under RDP/SSH constraints match Zero Trust and ransomware guidance to limit remote desktop services and constrain lateral movement [Rose et al., 2020; CISA, 2025a, b, c].
+Our empirical reductions in LMSK/BREK under RDP/SSH constraints match Zero Trust and ransomware guidance to limit remote desktop services and constrain lateral movement [7, 8, 9, 10].
 Relation to Zero Trust and segmentation.
-Zero Trust (NIST SP 800-207) calls for minimizing implicit trust and continuously evaluating access in context [Rose et al., 2020].
+Zero Trust (NIST SP 800-207) calls for minimizing implicit trust and continuously evaluating access in context [7].
 Our metrics quantify the residual ease of movement under candidate policies, complementing microsegmentation by highlighting edges whose removal yields the steepest probability-mass drop.
 Organization.
-Section 2 reviews the literature.
-Section 3 formalizes matrices, operators, and fixed-point computation.
-Section 4 defines LMSK/BREK.
-Section 5 presents properties and bounds.
-Section 6 provides examples and evidence.
+Section 2 reviews the literature; Section 3 develops the graph-theoretic underpinnings; Section 4 formalizes the matrices, operators ($\oplus$ and $\otimes$), and computation for the proposed model; Section 5 defines the metrics LMSK and BREK; Section 6 presents formal properties and bounds; Section 7 provides examples and evidence; and Section 8 concludes with discussion and future work.
 2.  Background and Related Work
@@
 vulnerability-driven traversal.
-Beyond exploit-based movement (e.g., T1210) [MITRE, 2025d], ATT&CK catalogs techniques enabling lateral movement without software flaws: Valid Accounts (T1078) and Remote Services (T1021, including RDP/SSH) describe abusing legitimate credentials to authenticate and operate on remote hosts [MITRE, 2025f, b, c, e].
+Beyond exploit-based movement (e.g., T1210) [4], ATT&CK catalogs techniques enabling lateral movement without software flaws: Valid Accounts (T1078) and Remote Services (T1021, including RDP/SSH) describe abusing legitimate credentials to authenticate and operate on remote hosts [6, 2, 3, 5].
 Our formulation treats interactive services as high-pivot edges when paired with account access, independent of CVE posture.
 Attack graphs and probabilistic security metrics.
-Classical attack graphs model preconditions/exploits and enumerate paths to goals [Phillips and Swiler, 1998; Sheyner et al., 2002].
+Classical attack graphs model preconditions/exploits and enumerate paths to goals [11, 12].
 Later work augments graphs with probabilities and costs.
-We deliberately do not encode per-CVE logic; instead we treat reachable services as coarse pivot opportunities whose probabilities compose multiplicatively with noisy-OR unions, trading granularity for deployability on routine network snapshots.
-Algebraic path problems and semiring views.
+We deliberately do not encode per-CVE logic; instead we treat reachable services as coarse pivot opportunities whose probabilities compose via a probabilistic path operator ($\otimes$) with probabilistic union ($\oplus$) across alternatives, trading granularity for deployability on routine network snapshots.
+Algebraic path problems and operator view.
 Many path computations admit uniform treatment via algebraic path frameworks replacing $(+,\times)$ by problem-specific $(\oplus,\otimes)$.
-We adopt multiplication for per-step success and probabilistic sum (noisy-OR) for unions, yielding a monotone iteration toward a least fixed point; this aligns with semiring-style closures, though our $\oplus$ is non-idempotent [Mohri, 2002; Sanmartín and Naranjo, 2022].
-Noisy-OR and causal independence.
-Our union operator is the standard noisy-OR from probabilistic graphical models, aggregating independent causal mechanisms with low elicitation burden [Heckerman, 1994].
+We adopt a probabilistic path operator $\otimes$ for per-step composition and a probabilistic union $\oplus$ for alternatives (noisy-OR), yielding a monotone iteration toward a least fixed point; this is analogous to algebraic path problems—we do not assume a semiring ($\oplus$ and $\otimes$ are non-idempotent, and distributivity holds only as an inequality) [13, 17].
+Probabilistic union (noisy-OR) and causal independence.
+Our probabilistic union operator corresponds to the standard noisy-OR from probabilistic graphical models, aggregating independent causal mechanisms with low elicitation burden [14].
 Generalizations (e.g., Noisy-MAX) clarify when such factorizations are appropriate.
 Epidemic spreading analogies.
-Multiplicative transmission along edges with union across parallel paths is mathematically analogous to network contagion; classic results connect topology to spread [Newman, 2002].
+Probabilistic composition along edges with probabilistic union across parallel paths is mathematically analogous to network contagion; classic results connect topology to spread [15].
 Our aim is risk measurement (expected reach/blast), not epidemic prediction, but the analogy explains the outsized effect of a small set of high-pivot edges.
 AD attack-path tooling.
-In enterprise Windows domains, BloodHound operationalizes graph-based lateral-movement analysis at scale [SpecterOps, 2025].
+In enterprise Windows domains, BloodHound operationalizes graph-based lateral-movement analysis at scale [16].
 We follow similar instincts (directed, typed edges) but compute a probability-weighted closure and provide system-level metrics for control evaluation.
 Zero Trust and ransomware guidance.
-NIST SP 800-207 sets the architectural principle; ransomware advisories emphasize limiting RDP/remote services to reduce both initial access and lateral movement [Rose et al., 2020; CISA, 2025a, b, c].
-2.1 Graph-theoretic underpinnings
-$k$-hop walks via adjacency powers.
-Let $A$ be the (directed) adjacency matrix of the network.
-Classical algebraic graph theory shows that $(A^{k})_{ij}$ equals the number of $i{\to}j$ walks of length kk.
-[^1] As usual, walks may repeat vertices; paths are walks with no repeated vertices.
-Our probabilistic closure works with walks.
-See, e.g., Godsil and Royle [2001]; Biggs [1993]; Diestel [2005].
-This formalizes the “$k$-hop” notion we use when bounding series $A+AW+\cdots+AW^{K-1}$.
-Boolean-OR reachability and transitive closure.
-Over the Boolean semiring $(\lor,\land)$, the transitive closure (reachability) can be computed by Boolean matrix powering and union,
-$R\;=\;A\lor A^{(2)}\lor\cdots\lor A^{(n)}\,$,
-or equivalently via Warshall’s dynamic program for transitive closure [Warshall, 1962].
-From a more algebraic perspective, Kleene-style closure over idempotent semirings and dioids gives a unifying view of reachability and shortest paths [Kozen, 1994; Gondran and Minoux, 2008].
-Noisy-OR style union for probabilistic graphs.
-When parallel, conditionally independent mechanisms can each enable a step (e.g., multiple services/credentials offering a pivot), the probability that at least one succeeds is the probabilistic sum
+NIST SP 800-207 sets the architectural principle; ransomware advisories emphasize limiting RDP/remote services to reduce both initial access and lateral movement [7, 8, 9, 10].
+3.  Graph-theoretic underpinnings
+3.1 Adjacency matrices and notation
+Let $A$ be the (directed) adjacency matrix of the network; $A_{ij}{=}1$ iff a policy-permitted, actively listening service on $j$ is reachable from $i$, and $A_{ii}{=}0$; later we use a pivot-potential matrix $W\!\in[0,1]^{n\times n}$ with $W\leq A$ (elementwise).
+[…]
+$A=\begin{bmatrix}0&1&1&0\\ 0&0&1&1\\ 0&0&0&1\\ 0&0&0&0\end{bmatrix}$
+Figure 1: Directed graph and its adjacency matrix AA.
+3.2 $K$-hop path reachability (Boolean closure)
+Let $A\in\{0,1\}^{n\times n}$ be the adjacency matrix ($A_{ij}{=}1$ iff there is an edge $i{\to}j$).
+Over the Boolean semiring $(\lor,\land)$, define the Boolean matrix product $(X\diamond Y)_{ij}:=\bigvee_{\ell=1}^{n}X_{i\ell}\land Y_{\ell j}$, Boolean powers $A^{[0]}:=I$, $A^{[1]}:=A$, and for $k\geq 1$, $A^{[k{+}1]}:=A^{[k]}\diamond A$.
+Claim (standard).
+For any $i,j$ and $k\geq 1$, $A^{[k]}_{ij}=1$ iff there exists a simple path from $i$ to $j$ of length at most $k$.
+(Proof sketch: if a $k$-step walk exists, prune cycles to obtain a simple path of length $\leq k$; conversely, a $k$-edge path witnesses $A^{[k]}_{ij}=1$.) See [27, §1.3].
+Hence the $K$-bounded path reachability is
+$R_{K}\;=\;I\lor A\lor A^{[2]}\lor\cdots\lor A^{[K]},$
+and the full transitive closure is $R_{n-1}$, since any walk of length $\geq n$ repeats a vertex and can be shortened without changing reachability.
+One may compute $R_{n-1}$ via Warshall’s dynamic program or view it as a Kleene-style closure over an idempotent semiring [20, 21, 22].
+[…]
+$A^{[2]}\;=\;A^{2}\ \text{(Boolean)}\;=\;\begin{bmatrix}0&0&1&1\\ 0&0&0&1\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix}$
+Figure 2: Length-2 paths. Boolean square $A^{[2]}$ (entry $(u,v)=1$ iff some $w$ gives $u\!\to\!w\!\to\!v$ under Boolean product).
+[…]
+$R^{(\leq 2)}\;=\;I\,\lor\,A\,\lor\,A^{[2]}\;=\;\begin{bmatrix}1&1&1&1\\ 0&1&1&1\\ 0&0&1&1\\ 0&0&0&1\end{bmatrix}$
+Figure 3: Boolean-OR reachability up to length 2. Combines self-reachability, 1-step, and 2-step paths.
+Computing reachability in practice (unweighted).
+We compute the $K$-bounded Boolean-OR reachability recursively using the Boolean product $\diamond$ and union $\lor$.
+  • Base case. $R_{1}\;=\;I\,\lor\,A.$
+  • Iteration. For $k\geq 1$,
+    $R_{k+1}\;=\;R_{k}\,\lor\,(R_{k}\diamond A),$
+    i.e., add all new (k+1)(k{+}1)-hop connections obtainable by one more Boolean step from the current frontier $R_{k}$.
+  • Final result / stopping. After K1K{-}1 iterations, return $R_{K}$. Equivalently, stop early at the fixed point if $R_{k+1}=R_{k}$ (no new reachability).
+Notes.
+This iterative union of Boolean powers is equivalent to Warshall’s transitive-closure algorithm and to the Kleene star on the Boolean semiring.
+In practice, bitset/word-parallel Boolean matrix multiplication for $\diamond$ makes each iteration fast; complexity is comparable to $O(n^{3})$ in dense form, with early stopping common in sparse graphs.
+3.3 Pivot potential
+Definition and example.
+Let $S_{ij}$ denote the enabling mechanisms (e.g., services/credentials) from $i$ to $j$.
+The pivot-potential matrix $W\in[0,1]^{n\times n}$ aggregates mechanisms on the same edge via probabilistic union, i.e.
+$W_{ij}\;=\;1-\prod_{m\in S_{ij}}\bigl(1-\pi(m)\bigr),\qquad W\leq A\text{ (elementwise)}.$
+[…]
+$W=\begin{bmatrix}0.00&0.80&0.30&0.00\\ 0.00&0.00&0.80&0.80\\ 0.00&0.00&0.00&0.30\\ 0.00&0.00&0.00&0.00\end{bmatrix}$
+Figure 4: Pivot potential matrix $W$ via noisy-OR aggregation over services on each edge.
+Probabilistic union (noisy-OR) across mechanisms.
+When parallel, conditionally independent mechanisms can each enable a step (e.g., multiple services/credentials offering a pivot), the probability that at least one succeeds is the probabilistic union
 $p_{\text{union}}\;=\;1-\prod_{\ell}(1-p_{\ell})\,$,
-the same aggregation used in network reliability for parallel components [Colbourn, 1987].
-In probabilistic graphical models, this is the standard noisy-OR (and its generalizations), which captures causal independence with one parameter per parent [Pearl, 1988; Heckerman, 1994; Diez, 1993].
-Our use of a noisy-OR union along with multiplicative step success places the model within these classical frameworks.
+the same aggregation used in network reliability for parallel components [23].
+In probabilistic graphical models, this is the standard noisy-OR (and its generalizations), which captures causal independence with one parameter per parent [24, 14, 25].
+In our framework we denote this union by $\oplus$ and pair it with a probabilistic path composition $\otimes$ for sequential steps, placing the model within classical algebraic-path/semiring views.
 4.  The Proposed Model
-Our model quantifies lateral movement exposure by representing the network as a directed graph and simulating the propagation of compromise probability.
-This requires two foundational matrices derived from network state and security policies.
-3.1 Foundational Matrices
-  1. 1.
-    The Adjacency Matrix ($A$): This is an $n\times n$ binary matrix representing the network’s underlying structure, where $n$ is the number of assets.
-An edge exists only if a service is both permitted by network policy and actively listening on the destination asset.
-[…]
-    We take $A_{ii}=0$ for all $i$ (no self-loops). Let $I$ denote the $n\times n$ identity matrix.
-  2. 2.
-    The Pivot Potential Matrix ($W$): This is an $n\times n$ matrix that encodes the inherent pivot potential of each edge.
-Instead of considering only the single most powerful service, the model treats each reachable service on an edge as an independent opportunity for an attacker.
-The entry $(W)_{ij}$ represents the total probability of a successful pivot from asset $i$ to $j$, calculated as the probabilistic union (or ”noisy-OR”) of the pivot potentials of all services available on that edge.
-    Let $S_{ij}$ be the set of all services reachable from asset $i$ to asset $j$. The pivot potential is calculated as:
-    $W_{ij}=1-\prod_{s\in S_{ij}}(1-\pi(s))$
-    Convention.
-$S_{ij}$ includes only services that are (i) allowed by policy and (ii) observed listening on $j$; hence $W_{ij}>0\Rightarrow A_{ij}=1$ (i.e., $W\leq A$ elementwise).
-In particular, if $S_{ij}=\varnothing$ then $W_{ij}=0$.
-3.2 Computing Reachability (Unweighted Method)
-Before detailing our probabilistic model, we first describe the standard, unweighted method for calculating K-hop reachability.
-This method answers the binary question: ”Does a path of length $K$ or less exist?” This is computed iteratively.
-Let $S_{k}$ be the matrix where $(S_{k})_{ij}=1$ if a path of at most $k$ hops exists.
-  • Base Case: The 1-hop reachability matrix is the adjacency matrix, $S_{1}=A$.
-  • Iteration: The reachability matrix for at most $k$ hops, $S_{k}$, is found by taking the previous state $S_{k-1}$ and adding any new paths formed by one additional hop:
-    $S_{k}=S_{k-1}\lor(S_{k-1}\cdot_{bool}A)$
-  • Final Result: The result is the matrix $S_{K}$ after K1K-1 iterations.
-While this method confirms path existence, it treats a high-risk path (e.g., via SSH) as identical to a low-risk one (e.g., via a database port), which our probabilistic model addresses.
-Let $S_{1}=I\lor A$ and $S_{k+1}=S_{k}\lor(S_{k}A)$; write $S_{K}$ for the $K$-hop boolean reachability matrix.
+Our model quantifies lateral-movement exposure on the directed graph using the matrices $A$ and $W$ defined in Sec.
+3, and computes $K$-hop compromise probabilities.
+We now describe the probabilistic operators and iteration.
 4.1 Computing Compromise Probability (Probabilistic Method)
@@
   1. 1.
-    The Probabilistic Composition Product ($\otimes$): This operator, replacing matrix multiplication, calculates the probability of new compromise paths forming by propagating existing compromise probabilities through the pivot potentials in $W$.
-For matrices $X$ and $Y$, the product $Z=X\otimes Y$ is:
-    $Z_{ij}=(X\otimes Y)_{ij}=1-\prod_{m=1}^{n}(1-X_{im}\times Y_{mj})$ (1)
+    The Probabilistic Composition Product ($\otimes$): This operator, replacing matrix multiplication, composes sequential steps and aggregates alternatives via the probabilistic union.
+For matrices $X$ and $Y$,
+    $Z_{ij}\;=\;(X\otimes Y)_{ij}\;=\;\bigoplus_{m=1}^{n}\bigl(X_{im}\,\otimes\,Y_{mj}\bigr)\;=\;1-\prod_{m=1}^{n}\Bigl(1-\bigl(X_{im}\,\otimes\,Y_{mj}\bigr)\Bigr).$ (1)
   2. 2.
-    The Probabilistic Union Operator ($\lor_{p}$): This operator, replacing matrix addition, correctly combines probabilities from independent sets of paths using the union rule.
-For matrices $X$ and $Y$, the union $Z=X\lor_{p}Y$ is:
-    $Z_{ij}=(X\lor_{p}Y)_{ij}=1-(1-X_{ij})(1-Y_{ij})$ (2)
+    The Probabilistic Union Operator ($\oplus$): This operator, replacing matrix addition, correctly combines probabilities from independent sets of paths using the union rule.
+For matrices $X$ and $Y$, the union $Z=X\oplus Y$ is:
+    $Z_{ij}=(X\oplus Y)_{ij}=1-(1-X_{ij})(1-Y_{ij})$ (2)
 Stopping rule. Iterate until $\|P_{k+1}-P_{k}\|_{\infty}<\varepsilon$ or $k=K$. Cost. Each iteration is $O(n^{3})$ and GPU-friendly (dense ops).
 Independence Assumption
-It is important to note that the probabilistic composition operator ($\otimes$) implicitly treats paths through different intermediate assets ($m$) as probabilistically independent events.
+It is important to note that the probabilistic composition ($\otimes$) treats the success events through different intermediates $m$ as conditionally independent.
 This is a standard assumption in many network models that makes the problem computationally tractable.
-We acknowledge that in some real-world scenarios, such as an attack using a single stolen credential, these paths may not be fully independent.
-Conservative bound. For correlated routes, we can replace $1-\prod_{m}(1-X_{im}Y_{mj})$ with $\max_{m}(X_{im}Y_{mj})$ (lower bound).
+We acknowledge that in some real-world scenarios, such as an attack using a single stolen credential, these events may not be fully independent.
+Conservative bound. For conditionally dependent routes, replace  $1-\prod_{m}\!\bigl(1-(X_{im}\otimes Y_{mj})\bigr)\,$ with $\max_{m}\bigl(X_{im}\otimes Y_{mj}\bigr)$ (lower bound).
 Using these operators, we compute $P_{K}$ via the following iterative algorithm:
 Step 1: Initialization (Base Case)
-The 1-hop compromise probability matrix, $P_{1}$, is the adjacency matrix $A$ with a diagonal of 1s, as an asset that is the starting point of an attack is by definition compromised.
+The 1-hop compromise probability matrix uses edge step-success on the first hop:
+$P_{1}\;=\;I\,\oplus\,W.$
+[…]
+$P_{1}\;=\;I\,\oplus\,W\;=\;\begin{bmatrix}1.00&0.80&0.30&0.00\\ 0.00&1.00&0.80&0.80\\ 0.00&0.00&1.00&0.30\\ 0.00&0.00&0.00&1.00\end{bmatrix}$
+Figure 5: Hop-1 baseline. $P_{1}=I\oplus W$ encodes self-reachability and probabilistic 1-hop step-success on edges.
 Step 2: Iteration
-For each hop from $k=1$ to K1K-1, we calculate the next state, $P_{k+1}$.
-The new state is the probabilistic union of the current state, $P_{k}$, with all the new compromise paths that can be formed by one additional hop.
-$P_{k+1}=P_{k}\lor_{p}(P_{k}\otimes W)$
+For each hop from $k=1$ to K1K-1, let $W_{1}:=W$ and define $W_{k+1}=W_{k}\otimes W$. Then compute
+$P_{k+1}\;=\;P_{k}\,\oplus\,W_{k+1}.$
+Note.
+Since $W_{1}=W$ already contains all 1-hop terms and $P_{k}$ accumulates $I\oplus W_{1}\oplus\cdots\oplus W_{k}$, the union with $P_{k}$ leaves shorter-hop mass unchanged and only adds longer-hop contributions.
+[…]
+$W_{2}\;=\;W_{1}\otimes W\;=\;\begin{bmatrix}0.00&0.00&0.64&0.67\\ 0.00&0.00&0.00&0.24\\ 0.00&0.00&0.00&0.00\\ 0.00&0.00&0.00&0.00\end{bmatrix}$
+Figure 6: Two-hop contribution. $W_{2}=W\otimes W$ shows exactly 2-hop mass.
 Step 3: Final Result
-After K1K-1 iterations, the resulting matrix is the final K-hop Compromise Probability Matrix, $P_{K}$.
-Hop-1 semantics. We treat the first hop as pure network reachability: $P_{1}=I\lor_{p}A$. Pivot potentials in $W$ affect only subsequent hops ($k\geq 2$) via $P_{k}\otimes W$.
+After K1K{-}1 iterations, return the cumulative matrix
+$P_{K}\;=\;P_{K-1}\oplus W_{K}\;=\;I\oplus W_{1}\oplus W_{2}\oplus\cdots\oplus W_{K},\quad\text{with }W_{1}:=W,\;W_{k+1}=W_{k}\otimes W.$
+[…]
+$P_{2}\;=\;P_{1}\oplus W_{2}\;=\;\begin{bmatrix}1.00&0.80&0.75&0.67\\ 0.00&1.00&0.80&0.85\\ 0.00&0.00&1.00&0.30\\ 0.00&0.00&0.00&1.00\end{bmatrix}$
+Figure 7: Up-to-2-hop compromise probability. Combine hop-1 with the 2-hop contribution: $P_{2}=P_{1}\oplus W_{2}$.
 5.  Quantifying Susceptibility and Blast Radius
@@
 Concept: This metric quantifies the overall ease of movement within the network.
-It answers the question: ”If we randomly select two different assets, what is the average probability that a successful lateral movement event can occur between them within $K$ hops?”
+It answers the question: ”If we randomly select a directed pair $i\!\to\!j$ with $i\neq j$, what is the average probability that compromise propagates within $K$ hops?”
 Formal Definition: The Lateral-Movement Susceptibility is the mean of all non-diagonal entries in the final compromise probability matrix, $P_{K}$.
 This focus on non-diagonal elements ensures the metric measures the probability of movement *between* distinct assets.
-$\text{LMS}_{K}=\frac{1}{n(n-1)}\sum_{i\neq j}(P_{K})_{ij}$ (3)
+$\mathrm{LMS}_{K}=\frac{1}{n(n-1)}\sum_{i\neq j}(P_{K})_{ij}$ (3)
 where $n$ is the total number of assets.
@@
 This metric is ideal for tracking the network’s security posture over time and measuring the effectiveness of segmentation controls.
+Worked example (K=2 on the running graph).
+From Fig. 7, the off-diagonal sum of $P_{2}$ is $4.17$; with $n=4$ there are n(n1)=12n(n{-}1)=12 off-diagonal entries, so
+$\mathrm{LMS}_{2}\;=\;\frac{4.17}{12}\;\approx\;0.35.$
 5.2 Blast-Radius Estimate (BREK)
-Concept: This metric quantifies the expected scope of compromise from an attack.
-It answers the business-critical question: ”Assuming an attacker can start at any asset, what percentage of our entire network do we expect to be compromised?”
+Concept: This metric quantifies the expected scope of compromise. It answers: ”If the attacker starts at a uniformly random asset, what percentage of the network do we expect to be compromised?”
 Formal Definition: The Blast-Radius Estimate is the mean of *all* entries in the final compromise probability matrix, $P_{K}$, expressed as a percentage.
 By including the diagonal elements (which are always 1), this metric correctly incorporates the initially compromised asset as part of the total compromise.
-$\text{BRE}_{K}=\left(\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}(P_{K})_{ij}\right)\times 100\%$ (4)
+$\mathrm{BRE}_{K}=\left(\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}(P_{K})_{ij}\right)\times 100\%$ (4)
 Interpretation: BREK provides a direct estimate of the potential scope of a breach.
@@
 This metric is designed to communicate exposure in clear, non-technical terms, helping to justify security investments and prioritize hardening efforts on the assets that contribute most to the overall blast radius.
+Worked example (K=2 on the running graph).
+From Fig. 7, the total sum of $P_{2}$ is $8.17$ over $n^{2}{=}16$ entries, hence
+$\mathrm{BRE}_{2}\;=\;\frac{8.17}{16}\times 100\%\;\approx\;51.06\%.$
 6.  Formal Properties of the Model
 Overview.
-We prove monotonicity, boundedness, and convergence; derive a $K$-hop series upper bound and a single-path product lower bound; recover boolean reachability in the unweighted limit; characterize the limit as the least fixed point $P^{*}=F(P)=P\oplus(P\otimes W)$; and show all results persist under service-class lifting.
+We prove monotonicity, boundedness, and convergence; derive a $K$-hop series upper bound and a single-path product lower bound; recover boolean reachability in the unweighted limit; characterize convergence by the fixed-point stopping condition $P_{K+1}=P_{K}$ (equivalently, $P_{K}=P_{K}\oplus W_{K+1}$); under our dependence policy cycles add no mass, so stabilization occurs by $K\leq n{-}1$; and show all results persist under service-class lifting.
 6.1 Monotonicity
@@
 $\forall i,j:(P_{k+1})_{ij}\geq(P_{k})_{ij}$
-This is guaranteed by the definition of the probabilistic union operator ($\lor_{p}$).
-This property ensures the model behaves logically, as the total probability of compromise only ever increases or remains constant as more paths are considered.
+This follows since $\oplus$ is monotone: with $W_{k+1}:=W_{k}\otimes W$, we have $P_{k+1}=P_{k}\oplus W_{k+1}$.
 6.2 Boundedness
 The values in the compromise probability matrix are always bounded within the interval $[0,1]$.
-The pivot potentials $\pi(s)$ are defined in $[0,1]$, and both the probabilistic composition ($\otimes$) and union ($\lor_{p}$) operators are closed over this interval.
+The pivot potentials $\pi(s)$ are defined in $[0,1]$, and both the probabilistic composition ($\otimes$) and union ($\oplus$) operators are closed over this interval.
 This property guarantees the stability of the model, ensuring that the outputs are always valid probabilities.
-5.3 Convergence
-The sequence of matrices $P_{1},P_{2},\dots$ is guaranteed to converge to a fixed-point matrix $P^{*}$, where further iterations no longer change the result.
-This is a direct consequence of the Monotonicity and Boundedness properties (a monotonically increasing sequence that is bounded from above must converge).
-Unlike unweighted reachability, which converges in at most n1n-1 steps, the probabilistic model may require more iterations for the probability values to become fully stable.
-This is because paths longer than n1n-1 can still contribute a small but non-zero probability to the total, incrementally refining the final result.
-In practice, the algorithm is governed by a stopping rule: iterate until the change between successive matrices falls below a predefined tolerance threshold, $\varepsilon$, or a maximum number of hops $K$ is reached.
-Formally, the iteration stops when $\|P_{k+1}-P_{k}\|_{\infty}<\varepsilon$ or $k=K$.
-By continuity of $\oplus$ and $\otimes$ on $[0,1]$, the limit $P^{*}$ satisfies $P^{*}=F(P^{*})$ with $F(P)=P\oplus(P\otimes W)$; thus $P^{*}$ is the least fixed point above $P_{1}$.
+6.3 Convergence (monotone bounded iteration)
+Initialize $P_{1}:=I\oplus W$ and update layer by layer
+$P_{k+1}\;:=\;P_{k}\;\oplus\;W_{k+1},\qquad W_{1}:=W,\;\;W_{t+1}:=W_{t}\otimes W\quad(t\geq 1).$
+Each update is monotone entrywise (both $\oplus$ and $\otimes$ are monotone), and all values lie in $[0,1]$; hence the sequence $\{P_{k}\}$ is monotone and bounded and therefore converges.
+We denote the limit by
+$P_{\infty}\;:=\;\lim_{K\to\infty}P_{K}.$
+Stopping test. In practice we stop when
+$P_{K+1}=P_{K}\quad\text{(within tolerance $\varepsilon$),}$
+i.e., the (K+1)(K{+}1)-hop layer $W_{K+1}$ adds no new mass.
+Thus, at stabilization the matrix satisfies the fixed-point equation $P_{K+1}=P_{K}$ implies $W_{K+1}=0$ wherever $P_{K}<1$ (which holds off-diagonal in our model), since $P_{K+1}=P_{K}\oplus W_{K+1}$.
+Dependence policy (edge sharing).
+We do not count as independent any walk that shares at least one edge with a shorter, already-accounted walk between the same endpoints; such a walk contributes no additional probability mass beyond the shorter walk.
+Remark (cycles and truncation).
+Under the edge-sharing policy above, any walk that contains a cycle can be shortened by excising the cycle and therefore shares edges with a shorter walk; it adds no new independent mass.
+Enforcing strict simple-path semantics would require per-walk history and is exponential in the worst case [30].
+In practice we use layered accumulation and stop early at a fixed point when $P_{K+1}=P_{K}$ (within tolerance), typically at $K\!\ll\!n$.
+As a hard upper bound, we truncate at K=n1K=n-1: any walk of length $\geq n$ necessarily repeats a vertex and thus contains a cycle [27, §1.3].
+6.4 Connection to Algebraic Path Problems
+We operate on $[0,1]$ with two primitives: a probabilistic union $\oplus$ and a probabilistic path composition $\otimes$.
+At the scalar level, $\oplus$ is the noisy-OR $a\oplus b=1-(1-a)(1-b)$ and $\otimes$ is an abstract operator for sequential steps (under conditional independence one may take $a\otimes b:=ab$).
+The matrix lifts are
+$(X\oplus Y)_{ij}=X_{ij}\oplus Y_{ij},\qquad(X\otimes Y)_{ij}=\bigoplus_{m=1}^{n}\bigl(X_{im}\otimes Y_{mj}\bigr).$
+These operators are monotone on $0,1]^{n\times n}$; $\oplus$ is commutative and associative (non-idempotent), and in our setting the super-distributive bounds hold
+$(X\otimes Y)\oplus(X\otimes Z)\;\geq\;X\otimes(Y\oplus Z),\qquad(Y\otimes X)\oplus(Z\otimes X)\;\geq\;(Y\oplus Z)\otimes X.$
+The layered update $P_{k+1}=P_{k}\oplus W_{k+1}$ thus computes an algebraic-path closure by monotone iteration; at stabilization the result $P$ satisfies the fixed-point equation $P=P\oplus(P\otimes W)$, analogous to algebraic path problems via least fixed-point iteration on a monotone operator [21, 22, 13, 17].
 6.5 Refinement of Binary Reachability
 This is the most critical property of our model. It establishes a formal link to the unweighted model and demonstrates the value of the probabilistic approach.
-Let $S_{K}$ be the final reachability matrix from the unweighted (Boolean) method, and $P_{K}$ be the final matrix from our probabilistic method. The following soundness relationship holds:
-$(P_{K})_{ij}>0\;\Rightarrow\;(S_{K})_{ij}=1$
+Let $R_{K}$ be the final reachability matrix from the unweighted (Boolean) method, and $P_{K}$ be the final matrix from our probabilistic method. The following soundness relationship holds:
+$(P_{K})_{ij}>0\;\Rightarrow\;(R_{K})_{ij}=1$
 Consequence (off-diagonal). For all $i\neq j$,
-$0\leq(P_{K})_{ij}\leq(S_{K})_{ij}.$
+$0\leq(P_{K})_{ij}\leq(R_{K})_{ij}.$
 Thus, Boolean reachability upper-bounds connectivity, while our probabilistic model provides a probability-weighted refinement (typically strictly smaller).
@@
 This directly demonstrates the model’s utility in reducing complexity and focusing on what matters.
-5.5 Connection to Algebraic Path Problems
-Our construction follows the algebraic-path template over $[0,1]$ using the probabilistic-sum (noisy-OR) as the “addition” $\oplus$ and ordinary multiplication as the “multiplication” $\otimes$:
-xy= $x\oplus y\;=\;1-(1-x)(1-y),\qquad x\otimes y\;$=\;xy.
-Matrix composition then uses these operations in the usual way:
-$(X\otimes Y)_{ij}\;=\;\bigoplus_{m=1}^{n}X_{im}\otimes Y_{mj}\;=\;1-\prod_{m=1}^{n}\bigl{(}1-X_{im}Y_{mj}\bigr{)},$
-which is exactly our probabilistic composition operator. (Here $\oplus$ corresponds to the $\lor_{p}$ operator used elsewhere.)
-Remark (semiring-like, non-distributive).
-The pair $(\oplus,\otimes)$ is associative, commutative, and monotone with identities 0 (for $\oplus$) and $1$ (for $\otimes$), but it is not distributive. For all $a,b,c\in[0,1]$,
-$(a\otimes b)\,\oplus\,(a\otimes c)\ \geq\ a\otimes(b\oplus c),\qquad(a\otimes c)\,\oplus\,(b\otimes c)\ \geq\ (a\oplus b)\otimes c,$
-with equality iff, respectively, $a\in\{0,1\}$ or $bc=0$, and $c\in\{0,1\}$ or $ab=0$. We do not rely on distributivity in our results.
-Fixed-point view.
-Let $F(P)=P\oplus(P\otimes W)$. Since $F$ is monotone on the complete lattice $0,1]^{n\times n}$ and bounded in $[0,1]$, iterating
-$P_{k+1}\;=\;P_{k}\oplus(P_{k}\otimes W)$
-computes the path-closure (least fixed point) up to either a prescribed hop limit $K$ or an $\varepsilon$-tolerance, grounding our iteration in standard algebraic-path theory.
-5.6 K-hop Series Bound for Lateral Propagation
-Define, for $K\geq 1$, the path-sum matrix
-$U_{K}\;=\;A\;+\;AW\;+\;AW^{2}\;+\;\cdots\;+\;AW^{K-1},$
-where products are the usual matrix products. The term $AW^{t}$ aggregates all paths of length t+1t{+}1 (one landing hop in $A$ followed by $t$ pivot steps in $W$).
-Proposition 1 (K-hop series upper bound).
-For all $K\geq 1$ and all $i\neq j$,
-$(P_{K})_{ij}\;\leq\;\min\!\bigl{(}1,\ (U_{K})_{ij}\bigr{)},\qquad\text{and}\qquad(P_{K})_{ii}=1.$
+6.6 Bounds for $K$-hop Lateral Propagation
+Upper bound by Boolean reachability (trivial).
+Let $R_{K}$ be the Boolean $K$-hop reachability matrix. Then for all $i\neq j$,
+$0\;\leq\;(P_{K})_{ij}\;\leq\;(R_{K})_{ij}\;\in\{0,1\},\qquad\text{and}\qquad(P_{K})_{ii}=1.$
+Reason. If $(P_{K})_{ij}>0$ then some $\leq K$ walk exists from $i$ to $j$, hence $(R_{K})_{ij}=1$. The diagonal equality follows from $P_{1}=I\oplus W$.
+Proposition 1 (single-path lower bound).
+Fix $K\geq L\geq 1$. Suppose there exists a directed path (no repeated vertices)
+$p:\ i=v_{0}\to v_{1}\to\cdots\to v_{L-1}\to v_{L}=j,$
+with step-successes $w_{r}:=W_{v_{r-1}v_{r}}\in[0,1]$ for $r=1,\dots,L$. Then
+$(P_{K})_{ij}\;\geq\;w_{1}\,\otimes\,w_{2}\,\otimes\,\cdots\,\otimes\,w_{L},$
+and under the scalar instantiation $a\otimes b:=ab$ this equals $\prod_{r=1}^{L}w_{r}$.
 Proof.
-Fix $i\neq j$.
-Let $\mathcal{P}_{\leq K}(i\!\to\!j)$ be the set of all directed paths from $i$ to $j$ of length at most $K$.
-By construction, $(P_{K})_{ij}$ is the probability of the union of the path events $\{E_{p}\}_{p\in\mathcal{P}_{\leq K}}$.
-By Boole’s union inequality, $\Pr(\bigcup_{p}E_{p})\leq\sum_{p}\Pr(E_{p})$.
-For a path of length $L\geq 1$ with vertices $i\to m_{1}\to\cdots\to m_{L-1}\to j$, the model assigns probability $A_{im_{1}}\prod_{r=1}^{L-1}W_{m_{r}m_{r+1}}$, so summing over all such paths of fixed length $L$ yields $(AW^{L-1})_{ij}$.
-Summing over $L=1,\dots,K$ gives $(U_{K})_{ij}$.
-Capping at $1$ preserves validity.
-The diagonal equality $(P_{K})_{ii}=1$ follows from the initialization $P_{1}=I\lor_{p}A$.
+  • By construction, $W_{1}=W$ and $W_{t+1}=W_{t}\otimes W$ expand all length-(t+1)(t{+}1) walks; hence the specific path $p$ contributes at least $w_{1}\otimes\cdots\otimes w_{L}$ to $(W_{L})_{ij}$.
+  • Since $P_{K}=I\oplus W_{1}\oplus\cdots\oplus W_{K}$ and $L\leq K$, monotonicity of $\oplus$ implies
+    $(P_{K})_{ij}\;\geq\;(W_{L})_{ij}\;\geq\;w_{1}\otimes\cdots\otimes w_{L}.$
 $\square$
-Corollary 1 (infinite-hops bound).
-If the spectral radius $\rho(W)<1$, then the Neumann series converges and
-$U_{\infty}\;:=\;\sum_{t=0}^{\infty}AW^{t}\;=\;A(I-W)^{-1},\qquad\text{hence}\qquad(P_{\infty})_{ij}\;\leq\;\min\!\bigl{(}1,\ (U_{\infty})_{ij}\bigr{)}.$
-Proof. Since $\rho(W)<1$, $\sum_{t\geq 0}W^{t}=(I-W)^{-1}$ converges. Apply Proposition 1 with $K\to\infty$ and linearity of matrix multiplication. $\square$
-Proposition 2 (single-path lower bound).
-Fix $K\geq L\geq 1$.
-Suppose there exists a directed path $p:\ i\!=\!m_{0}\to m_{1}\to\cdots\to m_{L-1}\to m_{L}\!=\!j$ in $A$ of length $L$, and let the pivot weights on its L1L-1 pivot steps be $w_{r}:=W_{m_{r}m_{r+1}}\in[0,1]$ for $r=1,\dots,L-1$.
-Then
-$(P_{K})_{ij}\;\geq\;\prod_{r=1}^{L-1}w_{r}.$
-Proof.
-$P_{K}(i,j)$ is the probability of a union over all $\leq K$-hop paths; it is at least the probability of any particular path event.
-Under the model’s hop-1 semantics ($A$ gives certainty on the first hop), the path $p$ succeeds with probability $\prod_{r=1}^{L-1}w_{r}$.
-$\square$
-Corollary 2 (uniform privileged pivots).
-Under the conditions of Proposition 2, if every pivot step on $p$ satisfies $w_{r}\geq p$ for some $p\in[0,1]$, then
-$(P_{K})_{ij}\;\geq\;p^{\,L-1}.$
-Proof. By monotonicity of products, $\prod_{r=1}^{L-1}w_{r}\geq\prod_{r=1}^{L-1}p=p^{L-1}$; combine with Proposition 2. $\square$
-5.7 Service-Class Lifting (Exact Land-then-Pivot Semantics)
-Not all footholds are equal: landing via an interactive remote-service (e.g., RDP/SSH) typically enables immediate lateral movement, whereas landing via application protocols (e.g., SQL/HTTP) rarely does without an exploit.
-To encode this service memory exactly, we lift states from nodes to (node, class) pairs.
-Classes and per-class adjacency.
+6.7 Service-Class Lifting via Node Splitting (Land–then–Pivot)
+The core idea of our advanced model is to distinguish how a foothold is landed (the service used to gain initial access to a node) from the service used for the next hop.
+These “land–then–pivot” semantics reflect a real-world observation: an attacker’s power to pivot from a server depends on how they landed on it in the first place.
+For example, if an attacker lands via a high-privilege service like SSH, the next hop succeeds with a high probability $\pi(\mathrm{SSH})$ as attacker needs sufficient privilege to make the next hop.
+Conversely, if the attacker lands via a low-privilege service like MSSQL, the next hop succeeds with a much lower probability $\pi(\mathrm{MSSQL})$.
+Thus, the hop success probability is a function of the landing service, not the outgoing service.
+This “memory” of the landing service is key to a more precise model.
+We retain a single adjacency $A\in\{0,1\}^{n\times n}$ on nodes $V$, and attach to each node $v\in V$ a small set of services $S(v)$ (e.g., $S(i)=\{\mathrm{SSH},\mathrm{MSSQL}\}$, $S(j)=\{\mathrm{SSH}\}$, …).
+Each service $s$ has a pivot rate $\pi(s)\in[0,1]$ that reflects the power an attacker has because they landed via $s$.
 […]
-$A\;=\;\bigvee_{c\in\mathcal{C}}A^{(c)},\quad\text{with }A_{ii}=0.$
-Let $\pi_{\mathrm{pivot}}(c)\in[0,1]$ denote the pivot probability conferred by a foothold acquired via class $c$.
-Lifted transition.
-A lifted state is $(j,c)$ meaning “at node $j$ with foothold class cc.” From $(j,c)$ one can move to $(\ell,c^{\prime})$ if and only if $A^{(c^{\prime})}_{j\ell}=1$, and the step succeeds with probability $\pi_{\mathrm{pivot}}(c)$.
-This defines the class-conditioned transition matrix $T\in[0,1]^{(n|\mathcal{C}|)\times(n|\mathcal{C}|)}$ with blocks
-$T_{(j,c)\to(\ell,c^{\prime})}\;=\;\pi_{\mathrm{pivot}}(c)\,\mathbf{1}\{A^{(c^{\prime})}_{j\ell}=1\}.$
-Initialization (hop 1: pure reach).
-From a start node $i$, the first hop ignores pivot and only records the landing class:
-$P^{\mathrm{lift}}_{1}\big{(}i\to(j,c)\big{)}\;=\;A^{(c)}_{ij},\qquad P^{\mathrm{node}}_{1}(i\to i)=1.$
-Iteration (hops $\geq 2$).
-For $k\geq 1$, propagate in the lifted space using the same operators as in Section 3.3:
-$P^{\mathrm{lift}}_{k+1}\;=\;P^{\mathrm{lift}}_{k}\;\lor_{p}\;\big{(}P^{\mathrm{lift}}_{k}\;\otimes\;T\big{)}.$
-Aggregation back to nodes.
-The $K$-hop node-to-node compromise probability is obtained by class-wise probabilistic union at the destination:
-$(P_{K})_{i\ell}\;=\;\bigl{(}\;\lor_{p}\;\{\,P^{\mathrm{lift}}_{K}(i\to(\ell,c))\ :\ c\in\mathcal{C}\,\}\bigr{)}.$
-All metrics (LMSK, BREK) continue to use $P_{K}$ exactly as defined here.
-Compatibility with the base model.
-If classes are collapsed by setting $W_{j\ell}=1-\prod_{c\in\mathcal{C}}\bigl{(}1-\pi_{\mathrm{pivot}}(c)\,\mathbf{1}\{A^{(c)}_{j\ell}=1\}\bigr{)}$, the lifted model reduces to the node-level model.
-Conversely, the lifted construction preserves hop-1 semantics and makes the dependence on the landing class explicit at every subsequent hop.
-Theorems unchanged.
-Proposition 1 (K-hop series upper bound) and Proposition 2 (single-path lower bound), and their corollaries, hold verbatim under this lifting.
-In Proposition 2 the per-step weights become the class-conditioned values $w_{r}=\pi_{\mathrm{pivot}}(c_{r})$ determined by the foothold class present at the source of step $r$.
-In Proposition 1, the path-sum bound is obtained by summing over class-typed paths; the resulting bound is still computable via ordinary matrix products by replacing $A$ with the block structure induced by $\{A^{(c)}\}$ and $W$ with $T$.
+$A=\begin{bmatrix}0&1&1&0\\ 0&0&1&1\\ 0&0&0&1\\ 0&0&0&0\end{bmatrix}$
+Figure 8: Directed graph and its adjacency matrix AA.
+$S(i){=}\{\mathrm{SSH},\mathrm{MSSQL}\},\quad S(j){=}\{\mathrm{SSH}\},\quad S(k){=}\{\mathrm{MSSQL}\},\quad S(\ell){=}\{\mathrm{SSH}\},$
+and rates
+$\pi(\mathrm{SSH}){=}0.8,\quad\pi(\mathrm{MSSQL}){=}0.3.$
+Lifted state space and transition.
+The lifted vertices are $(v,s)$ with $s\in S(v)$. We define a single lifted transition matrix
+$W^{\uparrow}_{(u,s_{\mathrm{in}})\to(v,s_{\mathrm{out}})}\;=\;\pi(s_{\mathrm{in}})\;\mathbf{1}\{A_{uv}=1,\ s_{\mathrm{out}}\in S(v)\}.$
+This ties the hop success to the landing service carried by the source state.
+In lifted order $(i,\mathrm{SSH}),(i,\mathrm{MSSQL}),(j,\mathrm{SSH}),(k,\mathrm{MSSQL}),(\ell,\mathrm{SSH})$,
+$(i,\mathrm{SSH})$$(i,\mathrm{MSSQL})$$(j,\mathrm{SSH})$$(k,\mathrm{MSSQL})$$(\ell,\mathrm{SSH})$0.80.80.80.80.30.30.30.30.80.80.80.80.30.3
+$W^{\uparrow}=\begin{bmatrix}0.0&0.0&0.8&0.8&0.0\\ 0.0&0.0&0.3&0.3&0.0\\ 0.0&0.0&0.0&0.8&0.8\\ 0.0&0.0&0.0&0.0&0.3\\ 0.0&0.0&0.0&0.0&0.0\end{bmatrix}$
+Figure 9: Lifted state graph $V^{\uparrow}=\{(v,s):s\in S(v)\}$ with transition weights equal to the pivot rate of the landing service.
+Initialization (lateral-only) and iteration in lifted space.
+Let $I^{\uparrow}$ be the identity on lifted states and $W^{\uparrow}$ the one-hop lifted operator. Define
+$W^{\uparrow}_{1}:=W^{\uparrow},\qquad W^{\uparrow}_{k+1}:=W^{\uparrow}_{k}\otimes W^{\uparrow}\quad(k\geq 1).$
+Accumulate the $K$-hop series using the usual operators:
+$P^{\uparrow}_{1}:=I^{\uparrow}\oplus W^{\uparrow}_{1},\qquad P^{\uparrow}_{k+1}:=P^{\uparrow}_{k}\oplus W^{\uparrow}_{k+1}\quad(k\geq 1).$
+$(i,\mathrm{SSH})$$(i,\mathrm{MSSQL})$$(j,\mathrm{SSH})$$(k,\mathrm{MSSQL})$$(\ell,\mathrm{SSH})$11111111110.80.80.80.80.30.30.30.30.80.80.80.80.30.3
+$P^{\uparrow}_{1}\;=\;I^{\uparrow}\oplus W^{\uparrow}\;=\;\begin{bmatrix}1.0&0.0&0.8&0.8&0.0\\ 0.0&1.0&0.3&0.3&0.0\\ 0.0&0.0&1.0&0.8&0.8\\ 0.0&0.0&0.0&1.0&0.3\\ 0.0&0.0&0.0&0.0&1.0\end{bmatrix}$
+Figure 10: Lifted one-hop reach $P^{\uparrow}_{1}$ (identity + one-hop).
+Collapse to nodes (single, consistent rule).
+After computing reachability in the lifted space, we return to node–to–node probabilities by a single noisy–OR over both the source and destination service labels:
+$(P_{K})_{uv}\;=\;\bigoplus_{s\in S(u)}\;\bigoplus_{t\in S(v)}\;P^{\uparrow}_{K}\!\big[(u,s)\to(v,t)\big].$
+We always apply this double union, for all $K$, and we perform it once at the end (i.e., no intermediate collapses).
+This corresponds to the modeling choice that an attacker can leverage any foothold service at the source node and, on a hop, can attempt all destination interfaces; independence across attempts is captured by the noisy–OR operator $\oplus$.
+Since $P^{\uparrow}_{K}$ includes $I^{\uparrow}$, node diagonals of $P_{K}$ are handled automatically.
+[…]
+$P_{1}\;=\;\begin{bmatrix}1.00&0.86&0.86&0.00\\ 0.00&1.00&0.80&0.80\\ 0.00&0.00&1.00&0.30\\ 0.00&0.00&0.00&1.00\end{bmatrix}$
+Figure 11: Node-level one-hop reach $P_{1}$ (identity + one-hop).
+$(i,\mathrm{SSH})$$(i,\mathrm{MSSQL})$$(j,\mathrm{SSH})$$(k,\mathrm{MSSQL})$$(\ell,\mathrm{SSH})$0.640.640.730.730.240.240.310.310.240.24
+$W^{\uparrow}_{2}\;=\;\begin{bmatrix}0.00&0.00&0.00&0.64&0.73\\ 0.00&0.00&0.00&0.24&0.31\\ 0.00&0.00&0.00&0.00&0.24\\ 0.00&0.00&0.00&0.00&0.00\\ 0.00&0.00&0.00&0.00&0.00\end{bmatrix}$
+Figure 12: Lifted exactly-two-hop transition $W^{\uparrow}_{2}=W^{\uparrow}\otimes W^{\uparrow}$.
+$(i,\mathrm{SSH})$$(i,\mathrm{MSSQL})$$(j,\mathrm{SSH})$$(k,\mathrm{MSSQL})$$(\ell,\mathrm{SSH})$11111111110.80.80.930.930.730.730.30.30.470.470.310.310.80.80.850.850.30.3
+$P^{\uparrow}_{2}\;=\;P^{\uparrow}_{1}\oplus W^{\uparrow}_{2}\;=\;\begin{bmatrix}1.00&0.00&0.80&0.93&0.73\\ 0.00&1.00&0.30&0.47&0.31\\ 0.00&0.00&1.00&0.80&0.85\\ 0.00&0.00&0.00&1.00&0.30\\ 0.00&0.00&0.00&0.00&1.00\end{bmatrix}$
+Figure 13: Lifted reach $P^{\uparrow}_{2}$ (identity + one-hop + two-hop union).
+[…]
+$P_{2}\;=\;\begin{bmatrix}1.00&0.86&0.96&0.81\\ 0.00&1.00&0.80&0.85\\ 0.00&0.00&1.00&0.30\\ 0.00&0.00&0.00&1.00\end{bmatrix}$
+Figure 14: Node-level reach $P_{2}$ (identity + union of all 1–2 hop paths via lifted space).
+What this shows.
+We use a single model: node splitting by service with land–then–pivot semantics.
+Hop success depends on the landing service carried by the source state, so all compositions are done in the lifted space and we collapse back to nodes only once via a noisy–OR over both source and destination service labels.
+This preserves landing–service memory across hops; any intermediate collapse to nodes would lose that memory and undercount reach.
+The resulting $P_{K}$ feeds directly into $\mathrm{LMS}_{K}$ and $\mathrm{BRE}_{K}$.
 6.8 Computational Cost
@@
 7.  Examples and Evidence
-In this section, we provide a practical demonstration of our model.
-We begin with a simple, illustrative example to clarify the computational steps.
-We then present a more complex scenario to provide evidence of the model’s utility in identifying non-obvious security risks that are missed by standard reachability analysis.
-6.1 Illustrative Example
-Consider a simple network of three assets: a Web Server (Asset 1), an Application Server (Asset 2), and a Database (Asset 3).
-For this example, the Adjacency Matrix ($A$) representing the reachable services, and the corresponding Pivot Potential Matrix ($W$) derived from those services, are defined as:
-$A=\begin{bmatrix}0&1&0\\ 1&0&1\\ 0&0&0\end{bmatrix}\quad W=\begin{bmatrix}0.0&0.2&0.0\\ 0.9&0.0&0.3\\ 0.0&0.0&0.0\end{bmatrix}$
-Using the iterative algorithm defined in Section 3, we compute the 2-hop compromise probability matrix, $P_{2}$:
-  • First, the 1-hop matrix $P_{1}$ is initialized from $A$.
-  • Then, we compute the 2-hop result: $P_{2}=P_{1}\lor_{p}(P_{1}\otimes W)$.
-This yields the final probability matrix:
-$P_{2}=\begin{bmatrix}1.0&1.0&0.3\\ 1.0&1.0&1.0\\ 0.0&0.0&1.0\end{bmatrix}$
-From this matrix, we calculate the final metrics for $K=2$:
-  • LMS2: The average of the 6 off-diagonal elements is $(1.0+0.3+1.0+1.0+0.0+0.0)/6=0.55$.
-  • BRE2: The average of all 9 elements is $(6.3)/9\times 100\%=70\%$.
-This simple example demonstrates how the high pivot potential of the connection from Asset 2 to 1 creates a high probability of compromise across the network, resulting in a high Blast-Radius Estimate.
 7.1 Evidence of Practical Utility
-We now provide evidence of the model’s value in a more complex scenario where the goal is to prioritize security controls.
-Consider a network with four assets: a User Workstation (Asset 1), a Web Server (Asset 2), a critical Database (Asset 3), and a Domain Controller (Asset 4).
-The reachable services are:
-  • Workstation (1) $\to$ Web Server (2) via a low-privilege web port, $\pi=0.1$.
-  • Web Server (2) $\to$ Database (3) via a standard SQL port, $\pi=0.2$.
-  • Workstation (1) $\to$ Domain Controller (4) via an RDP port left open for remote admin access, $\pi=0.9$.
-Analysis with the Unweighted Method: A standard reachability analysis (using the Boolean method) would identify two lateral movement paths originating from the Workstation: one to the Database (length 2) and one to the Domain Controller (length 1).
-Both are considered ”reachable.” A security team might mistakenly prioritize the longer path involving the Web Server, or treat both paths as equally important, failing to see the true concentration of exposure.
-Analysis with Our Probabilistic Model: Our model calculates the compromise probabilities. The 2-hop compromise matrix $P_{2}$ would show:
-  • The probability of compromising the Database from the Workstation, $(P_{2})_{13}$, would be low.
-The probability of the two-hop path is based on the composition of low-privilege pivots (approximately $0.1\times 0.2=0.02$).
-  • The probability of compromising the Domain Controller, $(P_{2})_{14}$, would be very high, dominated by the single high-privilege RDP hop (0.9).
-The BRE from the Workstation would be heavily weighted by the near-certain compromise of the Domain Controller.
+We reuse the running 4-asset graph for continuity: $i=$ User Workstation, $j=$ Web/App Server, $k=$ Mid-tier Service Host, $\ell=$ Domain Controller.
+Binary reachability would mark both $i{\to}j{\to}\ell$ (two hops) and any direct admin path as equally “reachable.” Our probabilistic model—via $P_{2}$ (Fig.
+7)—quantifies the asymmetry: high-privilege remote-admin edges dominate over multi-hop, low-privilege routes, which explains the uplift in BRE2 (and LMS2) and prioritizes the single most important control.
 7.2 Illustrative Field Example: Service-Class Contrast in an Enterprise Snapshot
@@
 Nevertheless, the observation is consistent with our model’s predictions and with practitioner guidance prioritizing constraints on interactive remote services.
-[Tyagi, 2025; MITRE, 2025b, f, c]
-Conclusion: This scenario provides clear evidence of the model’s value.
-While the unweighted method presents a noisy picture of undifferentiated risks, our probabilistic model immediately and quantitatively identifies the single RDP connection as the most significant threat to the network.
-It correctly filters the noise from the low-probability path and pinpoints the connection that must be prioritized for remediation, demonstrating its practical utility in guiding security decisions.
+[26, 2, 6, 3]
+8.  Conclusion and Future Work
+Conclusion.
+We presented a probabilistic, service-aware model for lateral-movement exposure that cleanly separates structure (adjacency $A$) from step success (pivot potential $W$), composes hops via $\otimes$, and aggregates paths via $\oplus$.
+The running example and field snapshot illustrate how the model prioritizes high-impact controls beyond binary reachability.
+Future work.
+  • Technique/dependency modeling: relax noisy-OR independence by conditioning on attack techniques and shared prerequisites (credentials, agent presence); explore dependency-aware unions.
+  • Modern environments: extend to containers, APIs, and service-mesh edges (ephemeral topologies, short-lived identities).
+  • Vuln/RCE coupling: optional integration of vulnerability evidence (e.g., exploitable RCEs) as multipliers on $W$ while keeping the baseline CVE-agnostic design.
+  • Business value weighting: incorporate asset criticality and blast impact (risk == probability $\times$ loss), with weighted BRE/LMS for decision support.
+  • Policy optimization: compute minimal-control sets (segmentation/identity constraints) that maximally reduce $\mathrm{BRE}_{K}$ under operational constraints.